The good news ... DCV is all fixed and working fine, and looks to be pretty well calibrated. Replacing U503 did the trick.
The bad news ... resistance is still not working properly, I've replaced Q401 and Q402 ... Q401 definitely helped, 402 was desperation on my part, I think it was ok. I've also eliminated U507 (LTC1050) by removing it and the meter actually runs perfectly fine, not sure if it's used outside of the test mode. So I've probably replaced a perfectly good part here.
Anyway, I now have the 100R, 1K and 10K ranges seeming to be reasonable, the 100K and up ranges read progressively lower ... if I'm being picky then I think the 10K range is slightly low also, so I have a suspicion about the resistor network.
I think I've figured out most of the constant current source ... U402 with the 10v reference creates a 500uA or 50uA constant source which it feeds through a 2K resistor (in U403), this creates a 1V or 100mV drop which the other half of U402 re-establishes across the other resistors in U403 (switched in by U401.) In the pack, along with the 2K, there are 1K, 10K, 100K and strangely 160K (which I'm confused by, it should be 200K, but this may be the issue.)
So that should give you 1mA, 100uA, 10uA, 5uA, and 500nA sources. I can measure the drops across each of the resistors so can see that the switching is all working and I get a pretty constant 100mV or 1V depending on the range.
The problem really seems to be the 160K ... if I measure a 100K resistor on the 100K range it reads 68K, so that would point to that resistor being wrong ... however on the 1M range it drops to 50K, and a 1M resistor measures as 280K.
The 160K resistor does actually drop fairly dramatically when I measure it, especially when not on my meter's low power mode, whereas the 1K, 2K and 10K are rock solid, and I'm wondering whether it suffered as a result of the previous failures.
Does this sound feasible? It seems like a custom part, so I may have to bodge something together to test my theory, and help appreciated.
For info, pin 1 is common, with 100k to pin 2, 200k to pin 3, 1k to pin 4, 10k to pin 5, and 2k to pin 6. Pins 7 and 8 appear not connected.
Lee.